Determine the amount of money earned in an interest bearing account at 4% over 5 years and compounded monthly, if the principal was $20,000
Answer #1
Step-1:Calculation of future value of principal
Future value of principal = Principal*(1+i)^n Where,
= 20000*(1+0.003333)^60 i = 4%/12 = 0.003333
= $ 24,419.45 n = 5*12 = 60
Step-2:Calculation of amount of money earned
Amount of money earned = Future Value - Investment amount
= $ 24,419.45 - $ 20,000.00
= $   4,419.45